OpenStudy (anonymous):

Find the distance between (3, 8) and (7, -11). Round to the nearest hundredth. quick medal

pooja195 (pooja195):

Use the distance formula :) \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

pooja195 (pooja195):

\[\huge d=\sqrt{(7-3)^2+(-11-8)^2}\] \[\huge d=\sqrt{(4)^2+(-19)^2}\] Ill let u finish it off :)

OpenStudy (anonymous):

19.416 so rounding to the nearest hundredth would be? not good at rounding

pooja195 (pooja195):

we 4^2=16 -19^2=361 \[\huge d=\sqrt{(16+361}\] \[\huge d=\sqrt{377}\] I assume you can leave it as that and if you wanted to round it you can round to the nearest hundredth which would be 19.42

OpenStudy (anonymous):

thanks

pooja195 (pooja195):

np